why do actinides have more than +3 oxidation states? why do actinides have greater complex forming tendency than lanthanides? why do actinides have the tendency to form oxocations?

  1. why do actinides have more than +3 oxidation states?
  2. why do actinides have greater complex forming tendency than lanthanides?
  3. why do actinides have the tendency to form oxocations?

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
actinoid exhibit larger oxidation states than lanthanoids beacuse of very small energy gap between 5f , 6d and 7f subshell . 
the energy will be decided on the besis on (n+l ) rule . 
5f = 5+3 =8 
6d = 6+2 = 8 
ALL THE THINGS COMES OUT TO BE SAME HERE , SO, THEY HAVE SAME ENERGY 
HOPE IT CLEARS 
ALL THE BEST ..